OpenStudy (loser66):

Find the total ways in which 6 objective type questions can be answered when each question has 5 choices. Please, hep

OpenStudy (loser66):

another one: There are 5 multiple choice questions in an exam. How many sequences of answers are possible if the first question has 3 choices, next two have 4 choices and the last two have 5 choices.

OpenStudy (loser66):

Among the problems, which quantity is greater?

OpenStudy (loser66):

@zepdrix

OpenStudy (mathmath333):

is \(\large 5^{6}\) the answer for 1st one

OpenStudy (loser66):

I think so!! Just want to make sure it is right.

OpenStudy (loser66):

\(\left(\begin{matrix}5\\1\end{matrix}\right)\) for each object and we have 6 objects, That gives us \(\left(\begin{matrix}5\\1\end{matrix}\right)^6= 5^6\)

OpenStudy (loser66):

How about the second one? Actually, I don't like statistics and probability. I took the course and got very few knowledge :(

OpenStudy (mathmath333):

for the second one i think it is = \(\large 3\times 4^{2}\times 5^{2}\)